OpenStudy (anonymous):

What is the simplified form of 3 over 4x plus 3 + 21 over 8 x squared minus 14x minus 15 ? 6 times the quantity 2 x plus 5 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 over the quantity 4 x plus 3 6 times the quantity x plus 1 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 times the quantity x plus 1 end quantity over the quantity 2x plus 5 end quantity times 4 x plus 3

OpenStudy (anonymous):

\[\frac{ 3 }{ 4x + 3 } + \frac{ 21 }{ 8^2 - 14x -15 }\]

OpenStudy (anonymous):

so far i factored the denominator for the second fraction and got: \[\frac{ 3 }{ 4x + 3 } +\frac{ 21 }{ \left( 4x +3 \right) \left( 2x - 5 \right) }\] but now im stuck I don't know what to do

OpenStudy (anonymous):

@iGreen do you mind helping me out?

OpenStudy (anonymous):

@Hero @campbell_st @Callisto @CaseyCarns @perl